OpenStudy (anonymous):

The human body can only withstand deceleration of -294 m/s^2 (30g's). People typically experience this amount of deceleration during car accidents. If a car going 30 m/s comes to a stop with this acceleration, how long does it take to stop?

OpenStudy (anonymous):

a=v/t so t=v/a t=30(m/s)/294.10(m/s^2) t=0.102 sec

OpenStudy (anonymous):

key words: with this acceleration; a=-294 m/s^2 a car going; v=30 m/s was this problem ignoring fluid resistance and force of friction between the car and the road?